# Scale-Dependent Radial Geometry and Metric Mismatch in Wasserstein Propagation for Reverse Diffusion

### 20 Mar 2026 v1

@@ 3 Reverse-diffusion interface and time-dependent radial geometry
This section sets up the reverse diffusion and isolates the scale-dependent geometry used in the theorem. Under weak log-concavity, Gaussian smoothing can stabilize large separations before short scales satisfy a Euclidean one-sided dissipativity bound. We formalize this through a radial profile of the learned reverse drift and the quantities extracted from its two ends.
@@
We use the Euclidean inner product and norm throughout, write $\mathcal{P}_{p}(\mathbb{R}^{d})$ for laws with finite $p$-th moment, and for probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ let $\Pi(\mu,\nu)$ denote the set of couplings of $\mu$ and $\nu$, i.e. the Borel probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$. We denote the $2$-Wasserstein distance by $W_{2}$, and parametrize reverse time by t:=Tst:=T-s.
@@ 3.1 Forward smoothing and reverse interface
where $S_{a}\mu$ is the pushforward of $\mu$ under $x\mapsto ax$. Thus forward time acts by a deterministic linear shrinkage followed by Gaussian convolution.
The corresponding reverse drifts are
@@ 3.2 Structural assumptions
Assumption 2 (Average score forcing in $L^{2}(p_{s})$).
@@
Assumption 3 (One-sided Lipschitz of the score error).
@@
Assumption 2 controls the size of the forcing term, whereas Assumption 3 is the theorem-level geometric condition on the score error that enters pairwise distance evolution through the one-sided Lipschitz $\ell(s)$. The latter is not vacuous and should not be read as a black-box global regularity requirement: Appendix G records standard sufficient conditions implying it.
@@ 3.3 A radial profile dictionary and admissible switches
What drives the proof is not a single Euclidean one-sided Lipschitz constant, but a family of radial lower envelopes $r\mapsto\underline{\kappa}_{s}(r)$, one for each noise level $s$. The key point is that one and the same formula already contains the near-field Euclidean obstruction and the far-field contraction reserve created by smoothing.
@@
Here $\alpha_{s}$ is the effective large-scale convexity of $p_{s}$, whereas $M_{s}$ measures the residual short-scale nonconvexity after smoothing.
@@
Figure 1: (a) Pointwise radial lower profile at a fixed noise level $s$: $\underline{\kappa}_{s}(r)$ starts at $\mathfrak{b}(s)$, crosses zero at $R(s)$, and approaches $\mathfrak{m}(s)$. (b) Window-level quantities used by the theorem: $\underline{m}(s_{0})=\inf_{u\in[s_{0},T]}\mathfrak{m}(u)$ and $\Gamma(s_{0})=\int_{0}^{s_{0}}\mathfrak{b}(u)\,\mathop{}\!\mathrm{d}u$. Thus $s_{0}$ is admissible exactly when $\underline{m}(s_{0})>0$. The right panel is schematic and does not represent the evolution of $R(s)$.
The left panel is pointwise in $s$, whereas the right panel records the stronger uniform-in-time requirement that will define admissible switches.
@@
For the standard VP schedule, admissibility can be certified by an explicit scalar threshold; see Proposition 8. While $M$ does not affect bare admissibility, it still enters the early metric through $M_{s}$.
The running infimum $\underline{m}(s_{0})$ appears because the early phase needs a positive far-field tail on the whole interval $s_{0},T]$, not only at one instantaneous noise level. Admissibility therefore guarantees the existence of a genuinely non-Euclidean contractive regime on the early window, but it does not require the near-field Euclidean obstruction to have vanished there. That remaining short-scale obstruction is exactly why an adapted switch metric is still needed before returning to $W_{2}$.
For each admissible switch $s_{0}$, the appendices construct an associated switch metric $W_{\varphi_{s_{0}}}$, an early damping rate $c(s_{0})$, and an affine-tail slope $\mathfrak{a}(s_{0})$.
@@ 4 A routing theorem from radial metric mismatch
Section 3 extracted the geometric inputs carried by the theorem: the radial lower envelopes $\underline{\kappa}_{s}$, the admissible switch set $\mathcal{S}_{\mathrm{adm}}$, and the late-window Euclidean load $\Gamma$. We now state the one-switch routing theorem.
@@
For any increasing concave function $\varphi:[0,\infty)\to[0,\infty)$ with $\varphi(0)=0$, define
@@ 4.1 Main theorem
Work under Assumptions 1, 2 and 3. Let $0=t_{0}<\cdots<t_{N}=T$ be a numerical grid, let $(Z_{t_{k}})_{k=0}^{N}$ be a Markovian discretization of the learned reverse SDE, and write $P_{k}:=P_{t_{k},t_{k+1}}$ and $\Psi_{k}$ for the exact and numerical one-step kernels. Assume that the ideal reverse SDE (12) admits a global solution $(\widetilde{X}_{t})_{0\leq t\leq T}$ on $[0,T]$ such that
@@
and that the one-step defects satisfy
@@
Fix $p>2$ and assume $\widehat{p}_{T}\in\mathcal{P}_{p}(\mathbb{R}^{d})$. Let $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ be a grid-aligned admissible switch, set $t_{s}:=T-s_{0}=t_{K}$, and let $\varphi_{s_{0}}$, $c(s_{0})$, and $\mathfrak{a}(s_{0})$ denote the associated switch metric, early damping rate, and affine-tail slope. Assume that the learned reverse SDE is strongly well posed on $[0,T]$, and that for every $u,v]\subseteq[0,t_{s}]$ and every initial coupling $\pi\in\Pi(\mu,\nu)$ with $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, there exists a coalescing reflection coupling on $[u,v]$ whose initial law is $\pi$, and that $\underline{g}(s_{0})>0,\overline{\mathfrak{b}}(s_{0})<\infty,\overline{G}(s_{0})<\infty,\Gamma(s_{0})<\infty.$ Let $\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})<\infty$ and satisfy
@@
Theorem 1 is a routing statement. On $0,t_{s}]$, the usable contraction is carried by the radial metric $W_{\varphi_{s_{0}}}$, not by Euclidean $W_{2}$. At $t_{s}$, the damped switch-time discrepancy is converted once back to $W_{2}$, which is exactly where the exponent $\theta_{p}=(p-2)/(2(p-1))$ enters; on $t_{s},T]$, the remaining error is then propagated in Euclidean geometry.
This perspective also clarifies the role of the exponent $\theta_{p}=(p-2)/(2(p-1))$. It enters only through the one-time conversion from $W_{\varphi_{s_{0}}}$ back to $W_{2}$; it plays no role in the early reflection-coupling contraction and no role in the late Euclidean propagation. Writing
for the actual switch-time discrepancy, the theorem controls this quantity through the explicit envelope provided by Lemma˜14 together with the admissible switch-time moment budget from Proposition˜2. The appendices make this interface concrete. In particular, Appendix˜D, especially Section˜D.2, shows that within the affine-tail concave class considered here and under only a $p$-moment budget at switch time, the exponent $\theta_{p}$ is structurally sharp: it is exactly the price of recovering quadratic Euclidean transport from a deliberately softer early metric. Moreover, Appendix˜E packages the right-hand side of (26) into the scalar objective $\mathcal{B}_{p}(s_{0})$ and proves the discrete optimal-switch corollary Corollary˜2; Section˜F.2 compares this route with the direct full-horizon Euclidean argument; and Appendix˜G records concrete sufficient conditions verifying the theorem-level hypotheses, including standard Lipschitz and linear-growth conditions for the reflection-coupling input, conservative proxy switches, and the explicit variance-preserving example in Proposition˜8.
@@ 5 Conclusion
This paper shows that, beyond globally regular regimes, reverse diffusion should not be summarized by a single Euclidean dissipativity constant. Under weak log-concavity, Gaussian smoothing can generate contraction first in the radial far field while short scales remain obstructed in the Euclidean metric underlying $W_{2}$. The far-field reserve and the near-field obstruction are thus not competing effects, but two manifestations of the same radial geometry. The main difficulty is therefore geometric rather than purely dynamical: contraction may already be present before it becomes visible in the Euclidean metric of the final objective.
Our contribution is to formulate this mismatch as a routing principle. The analysis should proceed in the first geometry that truly contracts, extract its gain on the early window, and return to $W_{2}$ only when that return becomes favorable. In our setting, the relevant early geometry is radial rather than Euclidean; reflection coupling is therefore not the source of this geometry, but the natural device for accessing its one-dimensional distance dynamics. The induced concave metric records the resulting early contraction, and a single switch transports the damped discrepancy back to the Euclidean metric of the terminal objective. The exponent $\theta_{p}$ is then the structural price of this return: within the affine-tail concave class and under a $p$-moment budget, it is sharp.
Our analysis is restricted to weakly log-concave targets and to theorem-level hypotheses. The present theory is also intrinsically one-switch: it does not yet cover multi-switch, strongly anisotropic, or more intricate multiscale geometries, nor does it aim at sampler-specific optimal complexity bounds. Extending the framework beyond one switch, beyond affine-tail concave switch metrics, or to settings with weaker regularity remains open.
@@
The appendix is organized as follows. Appendix˜A develops the reverse-diffusion background and the Gaussian-smoothing profile calculus that produces the quantities $\alpha_{s}$, $M_{s}$, $\mathfrak{m}(s)$, and $\mathfrak{b}(s)$. Appendix˜B transfers that profile to the learned drift and characterizes admissible switches. Appendix˜C proves early-window contraction in the switch metric via reflection coupling and packages the pre-switch initialization, discretization, and score-forcing contributions. Appendix˜D proves the one-time conversion from $W_{\varphi_{s_{0}}}$ back to $W_{2}$, including the explicit switch-time moment budget and the sharpness of the exponent $\theta_{p}$. Appendix˜E propagates the converted discrepancy through the late Euclidean window, while Appendix˜F combines all pieces into the main theorem, derives the optimized-switch corollary, and compares the one-switch route with the direct full-horizon Euclidean argument. Finally, Appendix˜G records operational supplements: proxy switches, verifiable sufficient conditions for the reflection-coupling input, and the variance-preserving worked example.
@@ Appendix B Profile transfer, Euclidean closure, and admissible switches
B.1 One profile, two endpoints
@@
B.2 Admissible switches
@@ Appendix C Early-phase damping in the switch metric
This section proves the early-window contraction used later. The proof has three steps: build a time-uniform switch-level lower envelope on $0,t_{s}]$; use reflection coupling only for two copies of the learned reverse SDE to obtain semigroup contraction in $W_{\varphi_{s_{0}}}$; and insert the three pre-switch error channels (initialization, discretization, and score forcing) at their entry times and propagate them to $t_{s}$ under that contraction.
Two couplings appear and should not be conflated. Reflection coupling is used only to prove the learned-semigroup contraction. When comparing the ideal and learned reverse SDEs over a short interval, we instead use synchronous coupling, so the common noise cancels and only the drift mismatch $\delta_{t}$ remains. Throughout this section, $P_{u,v}$ denotes the exact learned reverse semigroup on $[u,v]$, $P_{k}:=P_{t_{k},t_{k+1}}$ the exact learned one-step kernels, and $\Psi_{k}$ the numerical one-step kernels; we write
@@
Whenever Lemma 13 is invoked, we additionally use the theorem-level input stated there: strong well posedness of the learned reverse SDE on $0,t_{s}]$, existence of coalescing reflection couplings on every subinterval, and the finiteness conditions $\underline{g}(s_{0})>0$, $\overline{\mathfrak{b}}(s_{0})<\infty$, and $\overline{G}(s_{0})<\infty$.
Whenever Lemma 14 is invoked, we also use the theorem-level existence of a global ideal reverse solution
@@
C.1 From admissibility to a uniform early profile
@@ C.2 Reflection coupling and radial reduction
C.3 The switch metric as a radial Lyapunov gauge
@@ C.4 Early semigroup contraction
$q_{\varepsilon}(t)\leq e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\,q_{\varepsilon}(u)+\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\|F_{r}\|\,\mathop{}\!\mathrm{d}r+\sqrt{\varepsilon}\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}|\beta(r)|\,\mathop{}\!\mathrm{d}r.$

### 30 Mar 2026 v2

@@ 3 Reverse-diffusion interface and time-dependent radial geometry
Under weak log-concavity, Gaussian smoothing can stabilize large separations before short scales satisfy a Euclidean one-sided dissipativity bound. We encode this scale dependence through a radial profile of the learned reverse drift, whose two endpoints will later determine the residual Euclidean load and the far-field contraction reserve.
@@
We use the Euclidean inner product and norm throughout, write $\mathcal{P}_{p}(\mathbb{R}^{d})$ for laws with finite $p$-th moment, and denote by $\Pi(\mu,\nu)$ the set of couplings of probability measures $\mu,\nu$ on $\mathbb{R}^{d}$. We write $W_{2}$ for the $2$-Wasserstein distance and parametrize reverse time by t:=Tst:=T-s. Thus larger $s$ corresponds to earlier reverse time, whereas smaller $s$ corresponds to later reverse time.
@@ 3.1 Forward smoothing and reverse interface
where $S_{a}\mu$ is the pushforward of $\mu$ under $x\mapsto ax$. Thus forward time combines deterministic linear shrinkage with Gaussian convolution. The corresponding reverse drifts are
@@ 3.2 Structural assumptions
Assumption 2 (Score forcing in $L^{2}(p_{s})$).
@@
Assumption 3 (One-sided monotonicity control of the score error).
@@
Assumption 2 controls averaged score forcing, whereas Assumption 3 controls the pairwise geometric component of the score error that enters distance evolution through $\langle e_{s}(x)-e_{s}(y),x-y\rangle$. Thus Assumption 3 is a one-sided monotonicity condition rather than a full global regularity requirement. Global Lipschitz or symmetrized Jacobian bounds are stronger sufficient conditions implying it, but are not intrinsic to the theorem. The two assumptions are complementary rather than interchangeable. Concrete sufficient conditions, together with a comparison against global Lipschitz-type hypotheses, are given in Appendix G.
@@ 3.3 A radial profile dictionary and admissible switches
What drives the proof is not a single Euclidean one-sided Lipschitz constant, but a family of radial lower envelopes $r\mapsto\underline{\kappa}_{s}(r)$, one for each noise level $s$. The key point is that the same formula encodes both the near-field Euclidean obstruction and the far-field contraction reserve created by smoothing.
@@
Here $\alpha_{s}$ quantifies the effective large-scale convexity of $p_{s}$, while $M_{s}$ measures the residual short-scale nonconvexity after smoothing.
@@
Figure 1: (a) Pointwise radial lower profile at a fixed noise level $s$. (b) Window-level quantities used by the theorem: $\underline{m}(s_{0})=\inf_{u\in[s_{0},T]}\mathfrak{m}(u)$ and $\Gamma(s_{0})=\int_{0}^{s_{0}}\mathfrak{b}(u)\,\mathop{}\!\mathrm{d}u$. Thus $s_{0}$ is admissible exactly when $\underline{m}(s_{0})>0$. The right panel is schematic and does not represent the evolution of $R(s)$.
The left panel is pointwise in $s$, whereas the right panel illustrates the uniform-in-time far-field positivity that underlies admissible switches.
@@
The infimum $\underline{m}(s_{0})$ enforces a positive far-field margin throughout the entire early window $s_{0},T]$. Thus admissibility does not mean that direct Euclidean closure is already available; it certifies only that a contractive radial regime is present before the return to $W_{2}$. For each admissible switch $s_{0}$, the appendices construct the associated switch metric $W_{\varphi_{s_{0}}}$, its damping rate $c(s_{0})$, and its affine-tail slope $\mathfrak{a}(s_{0})$.
@@ 4 A routing theorem from radial metric mismatch
Section 3 produced the theorem input from a single object: the radial lower profile $\underline{\kappa}_{s}$. Its far-field endpoint yields the admissible switch set $\mathcal{S}_{\mathrm{adm}}$, and its near-field endpoint yields the Euclidean load $\Gamma$. The theorem below turns this one-profile, two-endpoint structure into a one-switch routing bound.
@@
For any increasing concave function $\varphi:[0,\infty)\to[0,\infty)$ with $\varphi(0)=0$, define the associated transport cost
@@ 4.1 Main theorem
Throughout this subsection, work under Assumptions 1, 2, and 3. Let $0=t_{0}<\cdots<t_{N}=T$ be a numerical grid, and let $(Z_{t_{k}})_{k=0}^{N}$ be a Markovian discretization of the learned reverse SDE initialized by
$Z_{t_{0}}\sim\widehat{p}_{T},$
with numerical one-step kernels $\Psi_{k}$, i.e.
$\operatorname{Law}(Z_{t_{k+1}}\mid Z_{t_{k}}=x)=\Psi_{k}(x,\cdot),\qquad x\in\mathbb{R}^{d},\quad k=0,\dots,N-1.$
Write $P_{k}:=P_{t_{k},t_{k+1}}$ for the exact learned reverse one-step kernels. Assume that the ideal reverse SDE (12) admits a global strong solution $(\widetilde{X}_{t})_{0\leq t\leq T}$ on $[0,T]$ such that
@@
Assume moreover that the one-step defects satisfy
@@
Fix $p>2$ and assume $\widehat{p}_{T}\in\mathcal{P}_{p}(\mathbb{R}^{d})$. Let $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ be a grid-aligned admissible switch, set $t_{s}:=T-s_{0}=t_{K}$, and let $\varphi_{s_{0}}$, $c(s_{0})$, and $\mathfrak{a}(s_{0})$ denote the associated switch metric, early damping rate, and affine-tail slope. Assume that the learned reverse SDE is strongly well posed on $[0,T]$. Assume moreover that, for every $u,v]\subseteq[0,t_{s}]$ and every initial coupling $\pi\in\Pi(\mu,\nu)$ with $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, there exists a coalescing reflection coupling on $[u,v]$ with initial law $\pi$. Suppose further that
$\underline{g}(s_{0})>0,\qquad\overline{\mathfrak{b}}(s_{0})<\infty,\qquad\overline{G}(s_{0})<\infty,\qquad\Gamma(s_{0})<\infty.$
Let $\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})<\infty$ satisfy
@@
Theorem 1 isolates a single geometric transition at $s_{0}$: before the interface, error is transported in the first geometry that actually contracts; after the interface, it is propagated in the Euclidean geometry in which the terminal discrepancy is measured. Thus $\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})$ is a damped pre-switch budget, and $\theta_{p}=(p-2)/(2(p-1))$ is the one-time price of returning from the softer switch metric to $W_{2}$. Equivalently, the theorem induces a scalar switch-selection objective on the admissible grid: minimize the right-hand side of (26) over admissible grid-aligned $s_{0}$; see Corollary 2. Appendix D.2 shows in Proposition˜3 that this conversion price cannot be improved uniformly within the affine-tail concave class under the same switch-moment assumption.
The advantage is entirely an early-window phenomenon. Under weak log-concavity, the radial profile $\underline{\kappa}_{s}$ can become contractive in the far field before direct Euclidean closure is available. Reflection coupling sees this profile through the radial separation process, and the adapted metric $W_{\varphi_{s_{0}}}$ converts that far-field reserve into genuine transport contraction while absorbing the near-field obstruction. For fixed $s_{0}$, the routed and direct bounds share the same late-window term, so the comparison reduces exactly to the early window. The direct route weights an error injected at noise level $u\in[s_{0},T]$ by the residual Euclidean factor $e^{\int_{s_{0}}^{u}\mathfrak{b}(v)\,\mathop{}\!\mathrm{d}v}$, whereas the routed route first damps the same input by $e^{-c(s_{0})(u-s_{0})}$ and returns to $W_{2}$ only once. Hence the routed bound can be strictly sharper on mismatch windows. Appendix F.2 makes the comparison exact, and Appendix G.4 makes it concrete in a closed-form variance-preserving example.
@@ 5 Conclusion
The main message of this paper is that, under weak log-concavity, end-to-end error in reverse diffusion should not be propagated in a geometry chosen a priori. The central issue is not whether $W_{2}$ is the final error metric, but whether it is the right geometry in which to transport error at every stage of the reverse trajectory. The resulting gain is therefore not only geometric but quantitative: direct full-horizon Euclidean propagation can be genuinely conservative on mismatch windows.
More broadly, our result suggests that reverse-diffusion theory under weak regularity should be organized around geometry-adaptive error propagation rather than uniform metric propagation. The present paper isolates this principle in the simplest setting where it can be proved cleanly: a single radial interface and theorem-level bounds. A natural next step is to understand whether the same principle persists beyond one switch, beyond radial geometries, under weaker verifiable conditions, and in complexity guarantees tailored to concrete samplers.
@@
The appendix follows the proof spine of the paper. Appendix˜A extracts the forward-time smoothing input and derives the quantities $\alpha_{s}$, $M_{s}$, $\mathfrak{m}(s)$, and $\mathfrak{b}(s)$. Appendix˜B transfers this profile to the learned reverse drift and characterizes admissible switches. Appendix˜C proves early-window contraction in the switch metric and packages the pre-switch initialization, discretization, and score-forcing terms. Appendix˜D proves the one-time return from $W_{\varphi_{s_{0}}}$ to $W_{2}$, including the sharpness of $\theta_{p}$. Appendix˜E propagates the converted discrepancy through the late Euclidean window, and Appendix˜F combines the pieces into the main theorem and the routed-versus-direct comparison. Appendix˜G records auxiliary comparison principles, conservative admissibility certificates, sufficient conditions for the early coupling input, and the explicit variance-preserving comparison result.
@@ Appendix B Profile transfer, Euclidean closure, and admissible switches
B.1 Endpoint structure of the lower radial profile
@@
B.2 Characterization of admissible switch times
@@ Appendix C Early-phase damping in the switch metric
This section proves the early-window contraction used later. The argument has three steps. First, we replace the family
$\bigl(r\mapsto\underline{\kappa}_{s}(r)\bigr)_{s\in[s_{0},T]}$
by a single switch-dependent lower envelope that is uniform on $0,t_{s}]$. Second, we use reflection coupling for two copies of the learned reverse SDE to obtain semigroup contraction in $W_{\varphi_{s_{0}}}$. Third, we insert the three pre-switch error channels—initialization, discretization, and score forcing—at their entry times and propagate them to $t_{s}$ under that contraction.
Two couplings are used, for different reasons. Reflection coupling is used only for the learned-semigroup contraction. By contrast, when we compare the ideal and learned reverse dynamics on a short interval, we use synchronous coupling so that the noise cancels and only the drift mismatch $\delta_{t}$ remains. Throughout this section, $P_{u,v}$ denotes the exact learned reverse semigroup on $[u,v]$, $P_{k}:=P_{t_{k},t_{k+1}}$ the exact learned one-step kernels, and $\Psi_{k}$ the numerical one-step kernels; we write
@@
From this point on, two theorem-level inputs are used repeatedly. Whenever Lemma 13 is invoked, we use strong well posedness of the learned reverse SDE on $0,t_{s}]$, existence of coalescing reflection couplings on every subinterval, and the finiteness conditions
$\underline{g}(s_{0})>0,\qquad\overline{\mathfrak{b}}(s_{0})<\infty,\qquad\overline{G}(s_{0})<\infty.$
Whenever Lemma 14 is invoked, we also use the existence of a global ideal reverse solution
@@
C.1 A uniform early-window profile from admissibility
@@ C.2 Reflection coupling and radial reduction
C.3 Construction and generator bound for the switch metric
@@ C.4 Early semigroup contraction
$q_{\varepsilon}(t)\leq e^{\int_{u}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\,q_{\varepsilon}(u)+\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\|F_{r}\|\,\mathop{}\!\mathrm{d}r+\sqrt{\varepsilon}\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}|\beta(r)|\,\mathop{}\!\mathrm{d}r.$